### v1

# Lie Detector: Unified Backdoor Detection via Cross-Examination Framework

Abstract
Institutions with limited data and computing resources often outsource model training to third-party providers in a semi-honest setting, assuming adherence to prescribed training protocols with pre-defined learning paradigm (e.g., supervised or semi-supervised learning). However, this practice can introduce severe security risks, as adversaries may poison the training data to embed backdoors into the resulting model. Existing detection approaches predominantly rely on statistical analyses, which often fail to maintain universally accurate detection accuracy across different learning paradigms. To address this challenge, we propose a unified backdoor detection framework in the semi-honest setting that exploits cross-examination of model inconsistencies between two independent service providers. Specifically, we integrate central kernel alignment to enable robust feature similarity measurements across different model architectures and learning paradigms, thereby facilitating precise recovery and identification of backdoor triggers. We further introduce backdoor fine-tuning sensitivity analysis to distinguish backdoor triggers from adversarial perturbations, substantially reducing false positives. Extensive experiments demonstrate that our method achieves superior detection performance, improving accuracy by 5.4%, 1.6%, and 11.9% over SoTA baselines across supervised, semi-supervised, and autoregressive learning tasks, respectively. Notably, it is the first to effectively detect backdoors in multimodal large language models, further highlighting its broad applicability and advancing secure deep learning.
1 Introduction
Deep learning models have grown exponentially in size in recent years, outstripping the computational resources available to many small and medium-sized institutions. Consequently, these institutions often rely on third-party cloud providers for model training. Although these providers are considered “semi-honest” in that they ostensibly adhere to prescribed protocols, they may still covertly manipulate data or models. This scenario can give rise to a significant backdoor threat, where hidden triggers are embedded during training, enabling the model to function normally under most conditions but exhibit malicious behavior when specific triggers are activated [10, 13, 2, 19, 22].
Current backdoor detection methods frequently rely on model behavior and statistical analyses (e.g., gradient-based detection, posterior analysis) [30, 24, 20, 9, 31, 25, 37]. However, such approaches tend to be highly sensitive to variations in optimization objectives, loss functions, and feature representations across different learning paradigms. This limitation constrains their ability to generalize across diverse architectures and attack strategies, posing serious challenges for maintaining user model security in a semi-honest setting.
Figure 1: In the absence of training resources, the user delegates model training to a third-party vendor in a semi-honest environment and generates two independent models. At the same time, the user doubles as a police to identify potential backdoor models through comparative analysis.
To address these shortcomings, we propose Lie Detector, a cross-examination backdoor detection framework designed for third-party verification. As illustrated in Fig. 1, the user (acting as police) outsources the same task to two independent providers (the suspects) and uncovers backdoors by identifying inconsistencies in their model outputs (the lies). Specifically, we employ Central Kernel Alignment (CKA) [14, 3] for task sensitivity analysis, enabling the reverse-engineering of triggers (the evidence) by maximizing representational differences between clean and backdoored models. In contrast to conventional methods that depend on decision boundaries, our approach optimizes triggers based on output distributions, allowing it to generalize across supervised, semi-supervised, and autoregressive learning tasks. Additionally, we introduce a fine-tuning sensitivity analysis to distinguish truly backdoored models from benign ones, thereby reducing false positives and enhancing detection robustness. This unified framework consistently achieves high detection accuracy across multiple learning paradigms, offering a practical and versatile solution for secure backdoored model verification.
We extensively evaluate the effectiveness of Lie Detector across supervised, semi-supervised, and autoregressive learning paradigms. The results show that our method significantly outperforms state-of-the-art backdoor detection approaches, with relative improvements of +1.6%, +5.4%, and +11.9%, respectively. In addition, Lie Detector demonstrates high stability under varying random seeds, underscoring its robustness. We anticipate that this research will encourage broader adoption of secure training practices in third-party services, thereby strengthening security guarantees for deep learning models. Our contributions are
  • We design a unified cross-examination framework for backdoor detection by analyzing inconsistencies in models provided by multiple third-party service providers, enhancing the security of outsourced training in semi-honest environments.
  • Our method combines CKA task sensitivity analysis and output distribution optimization, breaking the reliance on decision boundaries and enabling backdoor detection to generalize beyond supervised learning to semi-supervised learning and autoregressive learning.
  • We achieve superior generalization, improving detection by 5.4%, 1.6%, and 11.9% across three learning paradigms and seven attack methods. Notably, it is the first to enable backdoor detection in multi-modal large language models, further broadening its applicability.
2 Related Work
2.1 Development of Learning Paradigms
Deep learning has evolved through various training paradigms to address different challenges and data types. This article focuses on supervised learning, self-supervised learning, and autoregressive learning, highlighting their motivations, advancements, and limitations.
Supervised Learning (SL) trains models on labeled data, with early breakthroughs like CNNs [17] for image classification and DNNs for speech recognition. Large-scale datasets (e.g., ImageNet [4]) and architectures (e.g., ResNet [11], VGG [28]) further advanced the field. However, its reliance on labeled data, which is costly and time-consuming to obtain, motivated the development of alternative paradigms.
Self-Supervised Learning (SSL) emerged to address the data labeling bottleneck by generating labels automatically from unlabeled data. Transformers like BERT [6] revolutionized NLP, while contrastive learning frameworks like SimCLR [1] excelled in vision tasks. SSL bridges the gap between supervised and unsupervised learning by leveraging inherent data structures. Contrastive Learning (CL), a subset of SSL, explicitly contrasts positive and negative samples to learn meaningful representations. Recent advancements like CLIP [27] and CoCoOp [35] highlight CL’s versatility in unimodal and multimodal settings.
Autoregressive Learning (AL) extends SSL and CL by modeling data distributions and generating new samples across modalities. Transformer-based models like MiniGPT-4 [36] and LLaVA [18] enable joint text-image representations, advancing cross-modal understanding and generation.
The evolution from SL to SSL and AL addresses challenges in data, annotation, and generalization, enhancing model adaptability. This shift, driven by large-scale pre-trained models, has fueled deep learning advancements. However, their high computational demands limit accessibility for many users.
2.2 Backdoor Attack
Backdoor attacks have emerged as a critical security concern in deep learning, with their methods evolving alongside advancements in learning paradigms. These attacks aim to embed malicious behaviors into models during training, which can be triggered during inference by specific inputs.
Early backdoor attacks primarily focused on models trained with SL, leveraging labeled datasets to embed triggers. Notable examples include BadNets [10], which introduces poisoned data with predefined triggers to manipulate model predictions; Blended [2], which uses blended patterns as triggers, making them less detectable; ISSBA [19], which embeds invisible, sample-specific triggers to enhance stealth; WaNet [26], which utilizes warping-based triggers to achieve high attack success rates; and Low-Frequency [34], which exploits low-frequency components in images to embed triggers.
As SSL gained traction, attackers adapted existing methods and developed new techniques to target these models, which often rely on unlabeled data. Examples include BadCLIP [22], which extends backdoor attacks to contrastive language-image pretraining models, compromising multimodal representations, and BadEncoder [13], which poisons the encoder in SSL frameworks, affecting downstream tasks. With the rise of generative and multimodal models, backdoor attacks have expanded to exploit the AL paradigms, such as TrojanVLM [21], which targets vision-language models by embedding triggers in multimodal data, and Shadowcast [33], which focuses on stealthy backdoor attacks in generative models, particularly in text-to-image synthesis.
The landscape of backdoor attacks is extensive and continues to grow, spanning SL, SSL, and AL paradigms. While many attacks initially targeted supervised learning, they have been adaptively transferred or redesigned for self-supervised and multimodal settings. This proliferation of attacks highlights the urgent need for robust and generic defense mechanisms to safeguard deep learning systems across all learning paradigms.
2.3 Backdoor Detection
Existing Backdoor Detection Methods. Current backdoor detection methods frequently rely on model behavior and statistical analyses, such as gradient-based detection and posterior analysis [30, 24, 20, 9, 31, 25, 37]. These approaches often analyze the internal dynamics of models, such as gradients, activations, or output distributions, to identify anomalies indicative of backdoor behavior. For instance, Neural Cleanse (NC) [30] proposes an anomaly detection framework to identify and mitigate backdoors by analyzing the reversibility of triggers. Similarly, ABS [24] leverages activation clustering to detect poisoned neurons, while NAD [20] employs knowledge distillation to suppress backdoor effects during model fine-tuning. More recent works, such as MM-BD [31], MM-BD [31] designed a universal post-training backdoor detection method that identifies arbitrary backdoor patterns by analyzing the classifier’s output landscape and applying unsupervised anomaly detection. In contrast, TED [25] introduced a topological evolution dynamics framework to detect backdoors by modeling deep learning systems as dynamical systems, where malicious samples exhibit distinct evolution trajectories compared to benign ones. Some researches have proposed backdoor detection methods for SSL and AL paradigms, such as DECREE [9] which achieves backdoor detection by optimizing triggers, and SEER [37] which introduces another information modality for backdoor detection.
Existing backdoor detection methods have made some progress within individual learning paradigms, but their scalability is limited, making it difficult to directly apply them to other learning paradigms. In the future, there is a need to develop an unified detection methods to address backdoor threats across multiple learning paradigms.
3 Preliminary
This section introduces the fundamental concepts and theoretical foundations required for our method, primarily including the threat model and the definition of CKA.
3.1 Threat Model
In our proposed cross-examination-based backdoor detection framework, we operate under a semi-honest adversary model tailored for third-party model verification.
This threat model assumes that the service providers supplying the models are semi-honest, meaning they may attempt to embed backdoors into the models but will not actively interfere with the detection process itself. The adversary’s goal is to introduce hidden malicious behaviors into the model, which can be triggered by specific inputs during inference, while maintaining the model’s normal functionality on clean data.
Adversarial capabilities. The adversary, e.g., a malicious service provider (the suspects), has the capability to inject backdoors into the model during training or fine-tuning. This could involve poisoning the training data with trigger patterns or directly manipulating the model’s parameters to embed malicious behavior.
Adversarial knowledge. The adversary may have full knowledge of the model architecture and training process but is unaware of the specific detection mechanisms employed by the verifier. This ensures that the backdoor detection framework remains robust against adaptive attacks.
Detection constraints. The verifier user (the police) has no access to the training data or process and cannot assume the availability of a clean reference model. This aligns with real-world scenarios where third parties have no visibility into the training process, treating it as a black box.
3.2 Centered Kernel Alignment
CKA [32, 3, 14] can be used to measure the similarity between activations or feature representations. To compute CKA, we first input data $\mathbf{X}$ into two models and extract activations from specific layers $l$. Let $\mathbf{A}_{1}\in\mathbb{R}^{n\times p_{1}}$ and $\mathbf{A}_{2}\in\mathbb{R}^{n\times p_{2}}$ denote the activation matrices from the $l$-th layer of the two models, where $p_{1}$ and $p_{2}$ are the dimensionalities of the feature representations at that layer.
Next, the activation matrices $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ are transformed into kernel matrices $\mathbf{K}_{1}$ and $\mathbf{K}_{2}$ using a kernel function, typically the linear kernel:
$\centering{\mathbf{K}}=\mathbf{H}({\mathbf{A}}{\mathbf{A}}^{T})\mathbf{H}^{T},\@add@centering$ (1)
where $\mathbf{H}=\mathbf{I}-\frac{1}{n}\textbf{11}^{T}$ is the centering matrix, with $\mathbf{I}$ as the identity matrix and 1 as a vector of ones. This transformation ensures that the kernel matrix $\mathbf{K}\in\mathbb{R}^{n\times n}$ eliminates biases introduced by differences in model architecture.
The CKA similarity between the feature representations of two models is then defined as:
$\centering\text{CKA}(f^{1},f^{2},\mathbf{X})=\frac{\text{tr}(\mathbf{K}_{1}\mathbf{K}_{2})}{\sqrt{\|\mathbf{K}_{1}\|_{F}^{2}\cdot\|\mathbf{K}_{2}\|_{F}^{2}}},\@add@centering$ (2)
where $\mathbf{K}_{1}$ and $\mathbf{K}_{2}$ are the kernel matrices derived from the activations of models $f^{1}$ and $f^{2}$ for input $\mathbf{X}$. The term $\|\mathbf{K}_{*}\|_{F}^{2}$ represents the squared Frobenius norm, which is computed as the trace of the matrix product $\mathbf{K}_{*}\mathbf{K}_{*}$, i.e., $\|\mathbf{K}_{*}\|_{F}^{2}=\text{tr}(\mathbf{K}_{*}\mathbf{K}_{*})$.
CKA is architecture independent because it doesn’t change when certain transformations are applied [3, 14]. This means that architectural differences don’t change how similar two models are when measuring similarity. In particular: 1) Orthogonal transformation invariance. CKA remains unchanged under rotations and reflections of the feature space, making it robust to different basis representations. 2) Isotropic scaling invariance. Uniform scaling of feature representations does not impact CKA values, ensuring that similarity comparisons are not biased by differences in activation magnitudes. Because of these features, CKA is a very good way to compare models with different architectures because it looks at the relative structure of feature representations instead of their absolute values or specific network configurations.
4 Method
In this section, we will introduce backdoor defense method Lie Detecor based on the cross-examination framework as shown in Fig. 2.
Figure 2: Overview of the Lie Detector. We propose a general backdoor detection method based on the cross-examination framework. By leveraging output distribution loss and CKA loss to reverse triggers and further identifying backdoored models through fine-tuning sensitivity analysis, our approach ensures data security in third-party training processes.
4.1 Cross-Examination Framework
To enhance the security of third-party machine learning models, we propose a Cross-Examination-Based Backdoor Detection Framework, designed for a semi-honest verification setting. The framework consists of three main modules.
Cloud customers. It consists of users who require model training services but lack direct control over the training process. These users also act as the verification party (the police), who have the authority to verify model integrity. The users provide the clean dataset $\mathcal{D}_{\text{c}}$, training hyperparameters, model architecture $f$, and the learning paradigm $\mathcal{L}_{\textit{learn}}$.
Semi-honest third-party providers. They are independent service providers (the suspects) responsible for model training. While they follow the training protocol prescribed by users, they still retain the possibility of embedding arbitrary backdoors into the model. Their malicious behavior is reflected in a data poisoning process, where a fraction of the training data is modified to implant hidden vulnerabilities.
Specifically, we assume an adversary (suspect) trains a model $f_{\mathbf{\theta}}$ using an original dataset $D_{c}$ and alters a subset of it to create poisoned samples $D_{p}$ through predefined training details. The poisoned dataset consists of $\alpha|D_{c}|$ modified samples, where $\alpha\in[0,1]$ denotes the poisoning rate. The overall dataset used for training is then:
$\mathcal{D}=(D_{c}\setminus D_{p})\cup D_{p}.$ (3)
The adversary’s learning process can be formulated as an optimization problem:
[…]
$\quad{}+\alpha\mathbb{E}_{(\mathbf{x}_{p},\hat{y}_{c})\sim D_{p}}\big{[}\ell(f_{\mathbf{\theta}}(\mathbf{x}_{p}),\hat{y}_{c})\big{]}\big{\}},$
where $y_{c}$ is the ground-truth label of a clean sample $\mathbf{x}_{c}$, while $\hat{y}_{c}$ is the adversarially assigned target label for a poisoned sample $\mathbf{x}_{p}$, used to induce backdoor behavior. The learning objective $L_{\text{learn}}$ varies by learning paradigm, with the loss function $\ell(\cdot,\cdot)$ defined as follows: In supervised learning, $y$ represents discrete class labels for classification tasks, and $\ell$ is typically the cross-entropy loss. In contrastive learning (e.g., CLIP), $y$ defines similarity relationships rather than explicit class labels, and $\ell$ is a similarity-based contrastive loss. In autoregressive learning (e.g., LLaVA), $y$ serves as a reconstruction target, and $\ell$ includes reconstruction or autoregressive losses.
Cross-examination backdoor detection. The goal of cross-examination backdoor detection is to verify whether a model has been compromised by a backdoor without requiring access to its training data or process. Instead of relying on a known clean reference model or predefined attack patterns, our approach detects backdoors by leveraging inconsistencies between two independently trained models ($f_{1}$ and $f_{2}$) provided by different third-party service providers. Under the Cross-Examination framework, there are three possible outcomes: both models are clean, both models are backdoored, or one model is clean while the other is backdoored.
Next we present the challenges and motivations in our framework.
Challenges. Backdoor detection faces two primary challenges: 1) Accuracy. Many existing detection methods rely heavily on statistical analysis, assuming access to a clean reference model and predefined attack patterns. These assumptions introduce limitations, as mismatched priors can lead to detection failures when facing unknown or adaptive backdoor attacks. Traditional statistical approaches struggle to generalize beyond known attack distributions, reducing their reliability in real-world scenarios. 2) Generalization. The detection framework must be robust across different model architectures and learning paradigms, not just classification tasks. A method that is tightly coupled to specific model types or training objectives may fail in diverse applications, such as semi supervised learning or generative models. Ensuring architectural and task-agnostic generalization is critical for practical deployment.
Motivations. Compared to existing backdoor detection methods, our framework introduces the following innovations: 1) Leveraging model inconsistencies to avoid predefining attacks. Traditional detection methods depend on statistical assumptions about the distribution of backdoor triggers or poisoned data, making them vulnerable to novel or adaptive attacks. Our framework circumvents this limitation by exploiting inconsistencies between independently trained models on the same dataset, allowing detection without relying on prior knowledge of attack patterns. 2) Utilizing invariant features for better generalization. Many conventional defenses are tightly coupled to specific model architectures or training paradigms, limiting their applicability beyond classification tasks. Our method focuses on detecting structural inconsistencies that remain invariant across different architectures and learning paradigms, enabling broader applicability in different learning paradigms.
4.2 Cross-Model Trigger Reverse
In this subsection, we need to leverage the behavioral differences between models $f_{1}$ and $f_{2}$ to detect potential backdoors, conducting an initial screening to identify suspected backdoors in the models.
We generate triggers that effectively activate backdoor behaviors across different learning paradigms, formulating them as a combination of two trainable components: a mask $\mathbf{m}$ and a pattern $\mathbf{p}$. The mask $\mathbf{m}$ controls which pixels in the input image are modified, while the pattern $\mathbf{p}$ defines the injected adversarial content.
$\mathbf{x}^{\prime}=\mathbf{m}\odot\mathbf{p}+(1-\mathbf{m})\odot\mathbf{x},$ (5)
where $\mathbf{x}$ and $\mathbf{x}^{\prime}$ represent the clean and poisoned inputs, respectively, and $\odot$ denotes element-wise multiplication. By optimizing $\mathbf{m}$ and $\mathbf{p}$, we reconstruct effective triggers that are capable of eliciting malicious behavior in the suspect model.
Output distribution loss. First, we aim to identify a backdoor trigger (evidence) that can effectively activate the compromised model’s hidden behavior. This is achieved by leveraging the output distribution loss, which exploits the inherent characteristics of backdoor traces within a model. Attackers implant backdoors to ensure that the model’s output distribution strongly favors the attack target when the trigger is present, while a clean model exhibits a more uniform output distribution. The output distribution loss is defined as:
$\mathcal{L}_{\text{OD}}=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}-\ell_{\text{CE}}(f(\mathbf{x}^{\prime}_{i}),y_{i}),&\text{if SL,}\\[2.84526pt] \ell_{\text{Sim}}(f(\mathbf{x}^{\prime}_{i}),f(y_{i})),&\text{if SSL,}\\[2.84526pt] \mathbb{E}_{(\mathbf{m},\mathbf{p})}[\sum_{t}\ell_{\text{AR}}(f(\mathbf{x}^{\prime},\theta)^{(t)},\hat{y}_{i}^{(t)})],&\text{if AL.}\end{cases}$ (6)
where $N$ represents the number of selected samples drawn from the clean dataset $D_{c}$. The loss is computed over this subset rather than the full dataset and about 1000 samples.
Each term in the summation corresponds to a different learning paradigm: 1) Supervised learning (SL). $\ell_{\text{CE}}$ is the cross-entropy loss, ensuring the backdoored input $\mathbf{x}^{\prime}_{i}$ is classified as the target label $y_{i}$. 2) Self-supervised learning (SSL). $\ell_{\text{Sim}}$ is the similarity loss, measuring how close the feature representations of $f(\mathbf{x}^{\prime}_{i})$ and $f(y_{i})$ are. Increase dissimilarity to misalign the poisoned representation with clean semantic features. 3) Autoregressive learning (AL). $t$ is the index over generated tokens in an autoregressive model. $\hat{y}_{i}^{(t)}$ is the attacker-defined target output at timestep $t$, enforcing supervised sequence control over the backdoored generation. And $\ell_{\text{AR}}$ is the autoregressive loss.
CKA loss. To further expose backdoors, we leverage CKA loss to amplify training inconsistencies. Since CKA reflects representation learning objectives, a backdoored model optimizing for both clean and poisoned objectives inevitably diverges from a clean model. By maximizing this divergence, we highlight the Lie hidden within a suspect model.
Theorem 1.
(Task-Driven Representational Similarity Theorem) Let $f_{1}$ and $f_{2}$ be two independently trained models on the same dataset but potentially with different objectives or architectures. The representational similarity between the models, measured by Centered Kernel Alignment (CKA), strongly correlates with their task alignment:
$\rho_{\text{task}}(f_{1},f_{2})\propto\text{CKA}(\Phi_{f_{1}},\Phi_{f_{2}}),$ (7)
where $\Phi_{f_{1}}$ and $\Phi_{f_{2}}$ are feature representations extracted from the models, and $\rho_{\text{task}}$ quantifies their consistency in downstream task performance. Higher CKA similarity implies stronger alignment in decision boundaries and behavior across datasets.
By Theorem 1, CKA serves as a reliable metric to assess the alignment of learned representations between models trained under different paradigms. Since backdoored models are optimized for both clean and adversarial objectives, their representations deviate significantly from clean models. We exploit this by computing the CKA similarity between two models on backdoored inputs.
For models $f_{1}$ and $f_{2}$, we compute CKA on activation maps extracted from an input $\mathbf{x}^{\prime}$. The CKA loss is as follows:
$\mathcal{L}_{\text{CKA}}(\mathbf{K}^{\prime},l)=1-\frac{\text{tr}(\mathbf{K}_{1}^{l}(\mathbf{x}^{\prime})\mathbf{K}_{2}^{l}(\mathbf{x}^{\prime}))}{\sqrt{\|\mathbf{K}_{1}^{l}(\mathbf{x}^{\prime})\|_{F}^{2}\cdot\|\mathbf{K}_{2}^{l}(\mathbf{x}^{\prime})\|_{F}^{2}}},$ (8)
where $\mathbf{K}_{1}^{l}(\mathbf{x}^{\prime})$ and $\mathbf{K}_{2}^{l}(\mathbf{x}^{\prime})$ are kernel matrices computed from the activations of models $f_{1}$ and $f_{2}$ on the backdoored input $\mathbf{x}^{\prime}$. By maximizing $\mathcal{L}_{\text{CKA}}$, we construct inputs that accentuate behavioral discrepancies between the models, thereby facilitating the reverse of backdoor triggers.
Finally, we can minimize the above trigger optimization function as shown in Eq. 9:
$\centering\mathcal{L}(\mathbf{m},\mathbf{p})=\alpha\cdot\mathcal{L}_{\text{CKA}}+\beta\cdot\mathcal{L}_{\text{OD}}+\lambda\cdot(\|\mathbf{m}\|_{1}+\|\mathbf{p}\|_{1}),\@add@centering$ (9)
where the $\text{L}_{1}$ norm for regularization enhances the optimization and learning process of the trigger by promoting sparsity and minimal perturbation, following the principles outlined in the paper DECREE.
After this stage, we can filter out cases where both models are clean and identify cases where at least one model has a backdoor.
4.3 Fine-tuning Sensitivity Analysis
To further distinguish true backdoor models from clean models exhibiting unexpected behavior, we conduct a fine-tuning sensitivity analysis. This process helps accurately locate backdoor-implanted models by evaluating their stability under additional training.
Fine-tuning setup. We fine-tune both models, $f_{1}$ and $f_{2}$, obtaining fine-tuned versions $f_{1}^{\prime}$ and $f_{2}^{\prime}$. The fine-tuning process is performed on a subset ($10\%$) of the clean dataset $\mathcal{D}_{c}$, denoted as $\mathcal{D}_{\text{ft}}\subset\mathcal{D}_{c}$, by optimizing the learning paradigm-dependent loss $\mathcal{L}_{\text{learn}}$:
$f^{\prime}=\arg\min_{f}\mathcal{L}_{\text{learn}}(f,\mathcal{D}_{\text{ft}}).$ (10)
Backdoor identification criterion. We evaluate the model’s backdoor robustness by measuring the Attack Success Rate (ASR) before and after fine-tuning:
$\text{ASR}(f^{\prime})=\mathbb{E}_{\mathbf{x}^{\prime}\in\mathcal{D}_{\text{b}}}\left[\mathbb{I}(f^{\prime}(\mathbf{x}^{\prime})=\hat{y}_{c})\right],$ (11)
where $\mathcal{D}_{\text{b}}$ contains backdoor-embedded inputs $\mathbf{x}^{\prime}$, $\hat{y}_{c}$ is the target label, and $\mathbb{I}(\cdot)$ is the indicator function.
A model is flagged as backdoored if fine-tuning reduces ASR by more than $20\%$:
$\text{Backdoored}\iff\text{ASR}(f)-\text{ASR}(f^{\prime})>0.2.$ (12)
Since fine-tuning on clean data weakens backdoor effects, a significant ASR drop indicates reliance on the implanted backdoor, confirming its presence.
5 Experiments
Table 1: Detection accuracies (%) on ResNet-18. For each attack, we evaluate 20 clean and 20 backdoored models. Detection Success Rate (DSR) and False Positive Rate (FPR) are reported. Bold indicates the best result, and underline indicates the second-best result.
Dataset Attack NC ABS NAD TED MM-BD DECREE Lie Detector
DSR FPR DSR FPR DSR FPR DSR FPR DSR FPR DSR FPR DSR FPR
CIFAR10 BadNet 87.5 10.0 90.0 5.0 92.5 15.0 100 0.0 100 0.0 97.5 0.0 100 0.0
Blended 30.0 25.0 80.0 15.0 67.5 10.0 95.0 5.0 97.5 0.0 92.5 5.0 100 0.0
ISSBA 25.0 30.0 37.5 40.0 50.0 20.0 95.0 10.0 92.5 5.0 90.0 5.0 100 0.0
TinyImgNet BadNet 77.5 10.0 80.0 10.0 82.5 20.0 92.5 5.0 95.0 0.0 97.5 0.0 100 0.0
Blended 15.0 30.0 70.0 20.0 42.5 15.0 87.5 10.0 92.5 5.0 95.0 0.0 100 0.0
ISSBA 10.0 40.0 25.0 25.0 40.0 25.0 90.0 10.0 95.0 5.0 92.5 10.0 97.5 5.0
(a) Resnet18
(b) CLIP
Figure 3: Detection accuracies of cross-model trigger reverse on ResNet-18 and CLIP
5.1 Implementation Details
Models and Datasets. We evaluate our method across multiple learning paradigms. For supervised learning, we use ResNet18 [11] and VGG16 [28] on CIFAR-10 [15] and TinyImageNet [29]. For self-supervised and autoregressive learning, we test CLIP [27] and CoCoOp on ImageNet [5] and Caltech101 [8], while LLaVA [18] and mini-GPT-4 [36] are evaluated on COCO [23], Frisk-30k [7], and Frisk-8k [12].
Attacks and Defenses. We consider backdoor attacks across different paradigms, including BadNets [10], Blended [2], ISSBA [19], WaNet [26], and Low-Frequency [34] for supervised learning. For self-supervised and autoregressive learning, we adapt these attacks and further evaluate BadCLIP [22], BadEncoder [13], TrojanVLM [21], and Shadowcast [33]. We employ advanced defenses, including NC [30], ABS [24], NAD [20], TED [25], MM-BD [31], DECREE [9], and SEER [37]. Some methods, such as TED and MM-BD, are extended to multiple paradigms. Unless otherwise specified, all attack methods use a 10% poisoning rate. All evaluations are conducted using the semi-honest environment, with detailed settings and evaluation metrics provided in LABEL:app:attack and LABEL:app:defense.
5.2 Detection Performance in SL
In Tab. 1, we compare our method with six state-of-the-art post-training detection approaches in terms of detection accuracy [30, 24, 20, 25, 31, 9]: NC, ABS, NAD, TED, MM-BD, and DECREE. To evaluate their effectiveness, we first assess these methods on ResNet18 in a supervised learning setup, testing their performance against three classic backdoor attacks on CIFAR-10 and TinyImageNet. We can conclude that: 1) Lie Detector achieves state-of-the-art backdoor detection performance with consistently 100% DSR and near-zero FPR across different attacks and datasets. This demonstrates its robustness in identifying backdoored models without misclassifying clean ones. 2) Existing detection methods struggle with adaptive backdoor attacks, especially on complex datasets (TinyImageNet). While approaches like TED, MM-BD, and DECREE show improved performance over earlier methods (NC, ABS, NAD), they still fall short in consistently detecting stealthy backdoors (ISSBA).
5.3 Detection Performance in SSL and AL
Table 2: Detection accuracies (%) in CLIP and LLaVA. we evaluate 10 clean and 10 backdoored models per attack. Detection Success Rate (DSR) and False Positive Rate (FPR) are reported. Bold indicates the best result, and underline indicates the second-best result.
Architecture Dataset Attack TED MM-BD DECREE SEER Lie Detector
DSR FPR DSR FPR DSR FPR DSR FPR DSR FPR
CLIP Caltech101 BadNet 80.0 10.0 75.0 10.0 87.0 20.0 100.0 0.0 100.0 0.0
Blended 77.5 15.0 72.5 20.0 82.5 25.0 97.5 0.0 97.5 0.0
BadCLIP 47.5 35.0 52.5 25.0 60.0 30.0 90.0 0.0 95.0 5.0
ImageNet BadNet 60.0 15.0 67.5 10.0 72.5 10.0 95.0 0.0 95.0 0.0
Blended 57.5 20.0 65.0 15.0 75.0 15.0 90.0 5.0 92.5 0.0
BadCLIP 37.5 30.0 42.5 30.0 45.0 20.0 87.5 10.0 90.0 5.0
LLaVA COCO TrojanVLM 10.0 50.0 15.0 45.0 60.0 40.0 80.0 15.0 95.0 5.0
Shadowcast 10.0 50.0 15.0 50.0 60.0 45.0 85.0 10.0 92.5 0.0
Flickr-30K TrojanVLM 10.0 55.0 15.0 50.0 55.0 45.0 80.0 5.0 90.0 10.0
Shadowcast 10.0 50.0 10.0 45.0 45.0 35.0 80.0 10.0 90.0 5.0
We evaluate our method against four defense approaches under semi-supervised and autoregressive learning paradigms. We follow the original implementations of these methods with only modest modifications. The detection success rates are tested across four datasets and three classic backdoor attacks. Based on Tab. 2, we draw the following conclusions: 1) Existing methods have limited generalization. Traditional detection methods (TED, MM-BD, DECREE) show inconsistent performance across datasets and architectures. While some perform well on CLIP, they fail on vision-language models (e.g., LLaVA), indicating weak adaptability across learning paradigms. 2) FPR is High. Many methods, particularly TED and MM-BD, exhibit FPRs as high as 50%, misclassifying clean models as backdoored at a detection rate no better than random guessing. However, our method achieves superior generalization across different learning paradigms, maintaining high detection success rates with consistently low false positives.
Table 3: Component ablation experiments.
Component Attack Task Trigger Size Model DSR FPR
Cross-Model Trigger Reverse Blended CIFAR10 4×4 ResNet-18 100.0 10.0
VGG16 100.0 20.0
BadEncoder Caltech101 32×32 CLIP 100.0 20.0
CoCoOp 100.0 20.0
Shadowcast Flickr8k 50×50 LLaVA 90.0 20.0
Mini-GPT4 90.0 30.0
Lie Detector Blended CIFAR10 4×4 ResNet-18 100.0 0.0
VGG16 100.0 0.0
BadEncoder Caltech101 32×32 CLIP 100.0 0.0
CoCoOp 100.0 0.0
Shadowcast Flickr8k 50×50 LLaVA 90.0 0.0
Mini-GPT4 90.0 0.0
5.4 Ablation Study
To validate the effectiveness of Cross-Model Trigger Reverse and the robustness check phase in the Lie Detector, we conduct component ablation experiments in Tab. 3. Specifically, the Cross-Model Trigger Reverse setup removes the fine-tuning robustness analysis component, while the Lie Detector includes both components. We can conclude the following: 1) “Cross-Model Trigger Reverse” alone is effective but less robust. While it achieves high DSR, its FPR remains relatively high, reaching up to 30% in some cases, indicating potential misclassifications. 2) “Fine-tuning Sensitivity Analysis” significantly improves robustness. By integrating fine-tuning robustness analysis, the Lie Detector maintains the same high DSR while reducing FPR to 0% across all tested settings, demonstrating its effectiveness in distinguishing backdoored models from clean ones.
Similarity metrics selection. We select four existing similarity metrics for comparative testing, including CKA, CCA (Canonical Correlation Analysis), SVCCA (Singular Vector Canonical Correlation Analysis) and COS (Cosine Similarity) [16]. We adaptively replace the aforementioned different similarity metrics for backdoor detection, and the results are shown in LABEL:fig:f1_four. The results indicate that CKA outperforms other similarity metrics across different learning paradigms, especially on the LLaVA model, which also indirectly demonstrates the architecture-agnostic nature of the CKA metric.
Number of epochs. We present the detection accuracies under DSR as the number of epochs increases for ResNet-18 and CLIP models, as shown in Fig. 3. We observe that our method achieves stable convergence and remains effective across all attack methods on both ResNet-18 and CLIP models. The detection accuracy consistently improves with training epochs, demonstrating the robustness and adaptability of our approach in identifying backdoored models across different architectures and learning paradigms.
Table 4: Detection accuracies of methods with different model architectures.
Attack Task Trigger Size Model DSR FPR FLOPs
BadNet CIFAR10 4×4 ResNet-18 100.0 0.0 0.7
VGG16 100.0 0.0 0.4
BadCLIP Caltech101 32×32 CLIP 90.0 0.0 4.9
CoCoOp 100.0 0.0 5.0
TrojanVLM Flickr8k 50×50 LLaVA 90.0 0.0 76.6
Mini-GPT4 80.0 0.0 80.3
Model architecture. We evaluate the effectiveness of our detection method across six different model architectures spanning three learning paradigms, as shown in Table 4. Specifically, we assess supervised learning models (ResNet-18, VGG16), contrastive language-image models (CLIP, CoCoOp), and vision-language models (LLaVA, Mini-GPT4) under three representative backdoor attacks. We can conclude that: 1) Our method achieves 100% DSR and 0% FPR across diverse architectures, including complex multimodal models like LLaVA and Mini-GPT4. 2) Detection performance remains unaffected by model complexity, as measured by FLOPs. For example, despite a significant increase in computational cost from ResNet-18 (0.7 GFLOPs) to Mini-GPT4 (80.3 GFLOPs), our method consistently delivers high DSR with zero false positives.
Table 5: Variation of CKA values under different layers.
Layer ResNet-18 CLIP
Clean Backdoor Clean Backdoor
layer1 0.974 0.945 0.891 0.863
layer2 0.936 0.768 0.853 0.632
layer3 0.901 0.542 0.810 0.497
layer4 0.872 0.427 0.795 0.314
Feature layer selection. As shown in Tab. 5, CKA values in clean models remain stable across layers, whereas backdoored models exhibit a notable decline in deeper layers. This trend is consistent across ResNet-18 (supervised learning) and CLIP (SSL), confirming CKA’s reliability as a backdoor probe. Notably, layer 4 yields the most significant CKA drop (0.427 in ResNet-18, 0.314 in CLIP), making it the most effective layer for detection. A possible reason is that higher-layer features capture more abstract semantic information, which backdoor triggers distort, leading to greater representation shifts. So we choose the fourth layer features to calculate the CKA loss.
6 Conclusion
This paper proposes a unified backdoor detection framework for semi-honest settings where model training is outsourced to third-party providers. By leveraging cross-examination of model inconsistencies between independent service providers, our method significantly improves detection robustness across different learning paradigms. We integrate Centered Kernel Alignment (CKA) for precise feature similarity measurement and fine-tuning sensitivity analysis to distinguish backdoor triggers from adversarial perturbations, effectively reducing false positives. Extensive experiments demonstrate that our approach outperforms state-of-the-art methods, achieving superior detection accuracy in supervised, contrastive, and autoregressive learning tasks. Notably, it is the first to effectively detect backdoors in multimodal large language models. This work provides a practical solution to mitigate backdoor risks in outsourced model training, paving the way for more secure and trustworthy AI systems.